An arrow shaft when shot vertically losses it's 60% energy in 3 seconds. The maximum height that can be reached by arrows is?

Grade:9

1 Answers

Arun
5 years ago
Let the initial speed be u ,, so after 3 sec it speed will be 2/5u
so from first equation of motion 
v = u + at 
2/5u = u - 9.8×3
u = 9.8×3×5/3 = 49m/s
u = 49m/s
Now third equation of motion 
v^2=u^2+2as 
at maximum height , V = 0 
, So H = u^2/2g 
H = 49×49/2×9.8 
H = 122.5 m